# Missing Pieces: How Do Designs that Expose Uncertainty Longitudinally Impact Trust in AI Decision Aids? An In Situ Study of Gig Drivers

### v2

@@ 5.1. Metrics of Trust and Reliance
In the following sections, we use this notation to describe the statistical models that we fitted:
@@ 5.2.1. Effects on Trust
Our model (Table 1 in Section D.2) found that pre-interaction trust was significantly and positively correlated with their daily trust (pre_trust_score: $\beta=0.471,SE=0.119,p=0.00029$). Therefore, participants’ baseline trust persisted throughout the their interactions with the tool. Consistent with Figure 4, trust also increased significantly with each passing day (day: $\beta=0.130,SE=0.027,p<0.00001$). Also consistent with Figure 4, perceived accuracy was significantly and positively correlated with trust (estimate_accurate: $\beta=0.415,SE=0.167,p=0.01357$), but it had less of an impact on trust with each passing day of the user study (day:estimate_accurate: $\beta=-0.121,SE=0.037,p=0.00126$). This suggests that, by the end of the user study, participants’ trust was based less explicitly on perceived accuracy.
@@ 5.2.2. Effects on Reliance
Most participants indicated their desire to maintain their level of reliance on the schedule recommendation tool, corresponding to a reliance score of 0 ($\mu=0.038,\sigma^{2}=0.643$). Trust and reliance were not strongly correlated ($R^{2}=0.099$); some participants consistently expressed high reliance but also lower trust. The mean reliance score appeared to decrease over time, with the mean being lowest on Day 4, but we could discern no clear dependence on perceived accuracy (Figure 6 in Appendix D). To clarify the nature of these longitudinal effects, we fitted another LMM using lme4 and lmerTest. This model was similar to the model for trust, except the reliance score was the dependent variable, and we included the trust_score as an independent variable:
@@
Our model (Table 2 in Section D.2) did not find any significant effects for perceived accuracy (estimate_accurate) or pre-interaction trust (pre_trust_score). However, we did find two significant effects: a negative effect from the day, supporting our initial observation ($\beta=-0.220,SE=0.089,p=0.01444$), and a positive effect from the day:trust_score interaction ($\beta=0.058,SE=0.024,p=0.01640$). The latter suggests that reliance depended on perceived accuracy indirectly through the trust score. Participants who trusted the tool more were more likely to continue relying on it, with this effect strengthening over repeated interactions even as overall reliance weakened.
@@ 5.3.1. Pre-Interaction Trust
Next, we analysed the effects of our design conditions on trust and reliance, starting with pre-interaction trust. We found that the mean pre-interaction trust scores for Conditions (B)/(D)/(R)/(RH) were 3.338, 3.629, 4.183, and 3.367; Condition (B) had the lowest trust, and Condition (R) had the highest.
To verify these observations, we fitted an ordinary least squares (OLS) model for pre-interaction trust using the Python package statsmodels 0.14.2 (Seabold and Perktold, 2010), with the condition as an independent variable. Relative to Condition (B), the daily estimate Condition (D) did not significantly differ in pre-interaction trust (contrast (D)-(B): $\beta=0.290,SE=0.369,p=0.43188$); the ranged and hedged estimate Condition (RH) (contrast (RH)-(B): $\beta=0.028,SE=0.384,p=0.94139$) did not either. Yet, Condition (R) had significantly higher pre-interaction trust relative to Condition (B) (contrast (R)-(B): $\beta=0.845,SE=0.384,p=0.02764$) and Condition (RH) (contrast (RH)-(R): $\beta=-0.817,SE=0.391,p=0.03685$). Therefore, exposing uncertainty through ranged estimates initially improved participants’ trust. In Section D.3, we report on a similar analysis for post-interaction trust.
@@ 5.3.2. Longitudinal Trust and Reliance
Condition (D) decomposed estimated earnings on a daily basis, thus providing information irrelevant to uncertainty. Our models (Table 1 and Table 2 in Section D.2) indicate that this did not significantly improve trust (condition=(D): $\beta=-0.494,SE=0.341,p=0.15139$) or reliance (condition=(D): $\beta=0.028,SE=0.294,p=0.92461$) over the baseline Condition (B). While our trust model did find a significant, positive longitudinal effect for Condition (D) ($\beta=0.111,SE=0.053,p=0.03548$), Figure 5 suggests that this effect does not represent a practically significant trend.
@@
Condition (RH) added lexical hedging to the ranged earnings estimates in Condition (R). This condition was not significantly different from Condition (B) in terms of marginal or longitudinal effects on trust and reliance. On Day 6, participants in Condition (RH) reported significantly lower reliance than participants in Condition (R) (Figure 6; $\mu=0.714,-0.182;95\%\ \textrm{CIs}=(0.240,1.188)$, $(-0.649,0.286)$), the only such significant pairwise difference on a daily basis (see Table 3 in Appendix D). Combined with the pre-interaction trust of Condition (RH) being significantly lower than Condition (R) (Section 5.3.1), we conclude that the addition of lexical hedging in Condition (RH) reversed the gains in trust and reliance from Condition (R)’s range-based uncertainty.
@@ 6. Qualitative Analysis
Overall, 7 participants completed the exit interview after the longitudinal user study. Three of these were from Condition (D), one was from Condition (R), and three were from Condition (RH):
@@ 6.1. Motivations and Routines
Participants reported a diversity of driving motivations and routines, which impacted their perceptions of the schedule recommendation tool’s usefulness. P1 and P2 viewed gig driving as a primary source of income, and thus found more value in the tool’s earnings estimates: “[The tool was] definitely worthwhile, just because it gave me a number, a projection. […] They definitely motivate me to keep going the next day.” – P1 (D) Meanwhile, P3–P7 used gig driving to supplement other sources of income, but P4 still viewed his earning goals as important. Unlike other participants, P7 delivered with a bicycle in his spare time. He felt that his current commitment was insufficient to want to use the tool more: “If I take this job to a full time, take it seriously? I would [want to use it more].” – P7 (RH) Nevertheless, drivers found value in the tool regardless of their level of motivation. P1–P6 all reported challenges in estimating their potential earnings as a consequence of unpredictability in gig demand, pay, or location, or of gig platforms providing insufficient information. For instance: “Uber’s details that they offer to drivers through their interfaces are sorely lacking. So I’m grateful for the opportunity to interact with this tool.” – P6 (D)
@@ 6.2. RQ1: Perceptions of Accuracy
When evaluating the tool’s accuracy, participants weighed its recommendations against their own routines and intuitions. For P1, P3, P6, and P7, the tool was a reference for how they could perform in their existing routines, rather than something that would reshape their routines: “I still would’ve followed my routine. […] I was fortunate enough to at least have the tool make me a schedule based on the routine that I currently do.” – P3 (RH) However, P5 suggested that the tool could use a question-answering approach to nudge users into altering their routines, by first understanding their activity patterns and then suggesting modifications. When the tool was inaccurate, participants reacted in different ways. P1, P2, and P4 observed that instances of the tool being inaccurate decreased their desire to comply with the tool’s recommendations: “If I was making more than what it said, I would have done it more consistently on the schedule.” – P2 (D) P4’s reactions to inaccuracies were influenced by his expectations. He was motivated on one instance by the tool’s estimates exceeding his goals: “My target’s […] \$30 an hour. Because those levels were consistently below \$30, […] I wasn’t motivated to study it. But when I saw the 4 to 6 am, that kind of piqued my interest.” – P4 (RH) Maintaining consistent perceptions of accuracy over time was important for building trust in this context. P1, P2, and P5 indicated that the outcomes of their first one or two days of interaction impacted their willingness to follow the recommendations for the rest of the study. P4 and P5 indicated that their use of the tool would be strengthened longitudinally if they consistently perceived its predictions as being accurate: “Once I learned that it was accurate, and I had trust in it, and it was really helping, then I’d probably use it more and more.” – P4 (RH)
@@ 6.3. RQ2: Perceptions of Uncertainty
All participants in Conditions (R) and (RH) (P3–P5, P7) appreciated the presence of ranges. P5 compared the tool’s range to his own experiences: “I was always over the average. So, to me, I was kind of in my head using that as a low.” – P5 (R) For participants in the other conditions, P1 and P2 indicated that they would’ve preferred to have had ranges. However, P6 suggested that ranges may lead to disappointment when they are used for goal-setting: “If they earn less than [the higher number], then they probably might feel disappointed in the tool through no fault of its own, right? If you say \$12 to \$18, and it comes in at \$14, […] I could understand how folks might look at that as a let down.” – P6 (D) Thus, ranged-based uncertainty was useful for decision-making, but needed to be calibrated against expectations.
@@ 7.1. Key Findings and Implications
Interactivity could help to maintain trust over time. Losses of trust can also be mitigated retrospectively based on trust repair strategies (Pareek et al., 2024). Based on our qualitative findings, we hypothesise two mechanisms by which interactivity could help to maintain and repair trust. First, interactivity may enhance perceptions of control. The modes of interaction suggested by our participants, such as hourly breakdowns and question-answering, would assure users that the AI has the intent and agency to capture and learn from their preferences (Pareek et al., 2024). Second, interactivity could help users to better recall their experiences and decisions. Most interview participants could not recall whether their earnings significantly differed from the tool’s estimates until the interviewer probed further.
@@ 7.2. Limitations
Our work has two primary limitations. First, we cannot claim that the design of our tool was optimal for engendering trust. Our focus was on testing how designs for exposing uncertainty would impact trust and reliance. Thus, we attempted to isolate the effect of this design dimension by refining the tool through a formative study and pilot (Appendix B). Nevertheless, further improvements may have been possible. For instance, we could not provide retrospective breakdowns of participants’ earnings due to data availability limitations. Thus, design choices orthogonal to the exposure of uncertainty may have impacted participants’ trust and reliance.
Second, despite our best efforts, our sample of drivers was limited. These individuals were at least aware, if not active users, of the Gridwise app, and thus they may have been more focused on their outcomes than the general gig driver population. The trust of users in an AI decision aid is contingent upon their domain knowledge (Kim et al., 2023; Wysocki et al., 2023), and — as we demonstrate (Section 6.1) — the extent to which they integrate the decision aid into their existing routines. A future study aimed at a broader population of gig drivers could uncover additional insights by explicitly controlling for factors such as full-time status and driver tenure. We were also unable to reach participants who discontinued their participation. Future studies that follow up with such participants would be a valuable source of data on mechanisms of trust loss and repair in longitudinal settings.
@@ 7.3. Recommendations for Future Work
The paucity of similar longitudinal, in-situ studies in prior work is understandable given the logistical challenges we encountered. We stress the importance of observational studies to improve domain understanding as a basis for longitudinal interventional studies. Our pilot interviews helped us to design a tool that was task-aligned, which led participants to find value in it over repeated interactions. Furthermore, our study design aimed to increase perceived control while reducing user burden through flexibility in the scheduling of participation; customisability of the constraints on the tool; shorter survey instruments; and incremental compensation.
@@ Acknowledgements.
We thank Jenny T. Liang, Joshua Mateer, Jose M. del Alamo, Ian Yang, Chenyang Yang, Kaia Newman, Stephanie Milani, Niklas Kühl, and Nadia Nahar for their feedback and discussion on the study’s instruments, protocol, analysis, and manuscript. Also, we thank Linda Moreci for assistance with participant compensation, and Akshara Khare and Anagha Ravi Shankara for contributions to the recommendation tool’s implementation. This work was conducted in collaboration with Gridwise; we thank Kunal Prajapati, Karan Moudgil, Ryan Green, and Brian Finnamore for help with recruitment and data procurement. We were supported by a research grant from Mobility21 and the Tang Family Endowed Innovation Fund.
@@ A.4. Survey 4 — Post-Survey
    This item is an original question that prompts the participant to consider their interactions with other types of recommendation systems. It assesses the extent to which participants would appreciate increased exposure of uncertainty through ranged estimates in such systems.
@@ B.2. Methodology
To analyse these interviews, we used the same methodology as the longitudinal interviews (Section 4.3).
@@ B.3.1. Formative Questions
Since the interviews were semi-structured, the script below focuses on the guiding questions that we asked participants. The interviewer also probed participants further depending on their responses.
@@ B.3.2. Evaluative Questions
  • Do you feel that the recommended schedule gives you enough information to decide whether you would want to follow it?
@@ B.4. Participants
As with the longitudinal user studies (Section 4.1), participants were recruited from the user base of Gridwise in July 2023. Gridwise distributed recruitment messages to 500 users, but otherwise did not interact with participants. Recipients were sampled from users in the United States who had completed at least one gig in a platform linked to the Gridwise app over the week preceding recruitment. We recruited 4 interview participants:

### v4

@@ 5.1. Metrics of Trust and Reliance
We use this notation to describe our statistical models:
@@ 5.2.1. Effects on Trust
[…]
On the left, a line plot with seven points increases from 3.424 on Day 1 to 3.843 on Day 7. The general trend is upwards but not monotonic. 95% confidence intervals for each of the points are overlapping. On the right, a boxplot shows two boxes for each day: one red (perceived inaccurate) and one green (perceived accurate). For Day 1, the IQR of the red box is below the IQR of the green box. The pattern is less clear for subsequent days: the means of the red boxes are sometimes higher and sometimes lower, but the IQRs are all overlapping.
@@
Our model (Table 1 in Section D.1) found that participants’ pre-interaction trust was significantly and positively correlated with daily trust (pre_trust_score: $\beta=0.471,SE=0.119,p=0.00029$). Therefore, participants’ baseline trust persisted throughout their interactions with the tool. Consistent with Figure 4, their trust also increased significantly with each passing day (day: $\beta=0.130,SE=0.027,p<0.00001$). Also consistent with Figure 4, perceived accuracy was significantly and positively correlated with trust (estimate_accurate: $\beta=0.415,SE=0.167,p=0.01357$), but it had less of an impact on trust with each passing day of the user study (day:estimate_accurate: $\beta=-0.121,SE=0.037,p=0.00126$). This suggests that, by the end of the user study, participants’ trust was based less explicitly on perceived accuracy.
@@ 5.2.2. Effects on Reliance
Most participants indicated their desire to maintain their level of reliance on the schedule recommendation tool, corresponding to a reliance score of 0 ($\mu=0.038,\sigma^{2}=0.643$). Trust and reliance were not strongly correlated ($R^{2}=0.099$); some participants consistently expressed high reliance but also lower trust. The mean reliance score appeared to decrease over time, with the mean being lowest on Day 4, but we could discern no clear dependence on perceived accuracy (Figure 6 in Section D.2). To clarify the nature of these longitudinal effects, we fitted another LMM using lme4 and lmerTest. This model was similar to the model for trust, except the reliance score was the dependent variable, and we included the trust_score as an independent variable:
@@
Our model (Table 2 in Section D.1) did not find significant effects for either the tool’s perceived accuracy (estimate_accurate) or the pre_trust_score. However, two effects were significant: a negative effect from the day, supporting our initial observation ($\beta=-0.220,SE=0.089,p=0.01444$), and a positive effect from the day:trust_score interaction ($\beta=0.058,SE=0.024,p=0.01640$). The latter suggests that reliance depended on perceived accuracy indirectly through trust. Participants who trusted the tool more were more likely to continue relying on it; this effect strengthened over interactions even as overall reliance weakened.
@@ 5.3.1. Pre-Interaction Trust
Next, we analysed the effects of the tool’s design conditions on trust and reliance, starting with pre-interaction trust. Conditions (B)/(D)/(R)/(RH) had mean pre-interaction trust scores of 3.338, 3.629, 4.183, and 3.367; Condition (B) had the lowest, and Condition (R) had the highest.
To verify these initial observations, we used the Python package statsmodels 0.14.2 (Seabold and Perktold, 2010) to fit an ordinary least squares (OLS) model for pre-interaction trust, with the condition as an independent variable. Relative to Condition (B), the daily estimate Condition (D) did not significantly differ in pre-interaction trust (contrast (D)-(B): $\beta=0.290,SE=0.369,p=0.43188$); neither did the ranged and hedged estimate Condition (RH) (contrast (RH)-(B): $\beta=0.028,SE=0.384,p=0.94139$). Yet, Condition (R) had significantly higher pre-interaction trust relative to Condition (B) (contrast (R)-(B): $\beta=0.845,SE=0.384,p=0.02764$) and Condition (RH) (contrast (RH)-(R): $\beta=-0.817,SE=0.391,p=0.03685$). Therefore, exposing uncertainty through range-based estimates initially improved participants’ trust. In Section D.3, we report on a similar analysis for post-interaction trust.
@@ 5.3.2. Longitudinal Trust and Reliance
[…]
Two line plots with five lines; the lines in the left plot each have seven points representing mean trust scores on Days 1–7, while the lines in the right plot each have six points representing mean reliance scores on Days 1–6. The lines represent the overall mean and Conditions (B), (D), (R), and (RH). Between the lines, Condition (R) is consistently higher than all others; the others are all close to each other, but Condition (RH) tends to have marginally lower trust and Condition (B) marginally lower reliance.
@@
Condition (D) decomposed estimated earnings on a daily basis, thus providing information irrelevant to uncertainty. Our models (Tables 1 and 2 in Section D.1) indicate that this did not significantly improve trust (condition=(D): $\beta=-0.494,SE=0.341,p=0.15139$) or reliance (condition=(D): $\beta=0.028,SE=0.294,p=0.92461$) over the base Condition (B). While our trust model did find a significant, positive longitudinal effect for Condition (D) ($\beta=0.111,SE=0.053,p=0.03548$), Figure 5 suggests that this effect does not represent a practically significant trend.
@@
Condition (RH) added lexical hedging to the range-based earnings estimates in Condition (R). This condition was not significantly different from Condition (B) in marginal or longitudinal effects on trust and reliance. On Day 6, participants in Condition (RH) reported significantly lower reliance than participants in Condition (R) (Figure 6; $\mu=0.714,-0.182;95\%\ \textrm{CIs}=(0.240,1.188)$, $(-0.649,0.286)$), the only such significant pairwise difference on a daily basis (see Table 3 in Section D.2). Combined with the pre-interaction trust of Condition (RH) being significantly lower than Condition (R) (Section 5.3.1), we conclude that the addition of lexical hedging in Condition (RH) reversed the gains in trust and reliance from Condition (R)’s range-based uncertainty.
@@ 6. Qualitative Analysis
Overall, 7 participants completed the exit interview after they completed all seven days of the longitudinal user study. Three of these were from Condition (D), one was from Condition (R), and three were from Condition (RH):
@@ 6.1. Motivations and Routines
Participants reported a diversity of motivations and routines for gig driving, which impacted their perceptions of the schedule recommendation tool’s usefulness. P1 and P2 viewed gig driving as a primary source of income, and thus found more value in the tool’s earnings estimates: “[The tool was] definitely worthwhile, just because it gave me a number, a projection. […] They definitely motivate me to keep going the next day.” – P1 (D) Meanwhile, P3–P7 used gig driving to supplement other sources of income, but P4 still viewed his earning goals as important. Unlike other participants, P7 delivered with a bicycle in his spare time. He felt that his current commitment was insufficient to want to use the tool more: “If I take this job to a full time, take it seriously? I would [want to use it more].” – P7 (RH) Nevertheless, drivers found value in the tool regardless of their level of motivation. P1–P6 all reported challenges in estimating their potential earnings as a consequence of unpredictability in gig demand, pay, or location, or of gig platforms providing insufficient information. For instance: “Uber’s details that they offer to drivers through their interfaces are sorely lacking. So I’m grateful for the opportunity to interact with this tool.” – P6 (D)
@@ 6.2. RQ1: Perceptions of Accuracy
When evaluating the tool’s accuracy, participants weighed its recommendations against their own routines and intuitions. For P1, P3, P6, and P7, the tool was a reference for how well they could perform in their existing routines, rather than something to reshape their routines: “I still would’ve followed my routine. […] I was fortunate enough to at least have the tool make me a schedule based on the routine that I currently do.” – P3 (RH) However, P5 suggested that the tool could use a question-answering approach to nudge users into altering their routines, by first understanding their activity patterns and then suggesting modifications. When the tool was inaccurate, participants reacted in different ways. P1, P2, and P4 observed that instances of the tool being inaccurate decreased their desire to comply with the tool’s recommendations: “If I was making more than what it said, I would have done it more consistently on the schedule.” – P2 (D) P4’s reactions to inaccuracies were influenced by his expectations. He was motivated on one instance by the tool’s estimates exceeding his goals: “My target’s […] \$30 an hour. Because those [estimated earnings] were consistently below \$30, […] I wasn’t motivated to study it. But when I saw the 4 to 6 am, that kind of piqued my interest.” – P4 (RH) Maintaining consistent perceptions of accuracy over time was important for building trust in this context. P1, P2, and P5 indicated that the outcomes of their first one or two days of interaction impacted their willingness to follow the recommendations for the rest of the study. P4 and P5 indicated that their use of the tool would be strengthened longitudinally if they consistently perceived its predictions as being accurate: “Once I learned that it was accurate, and I had trust in it, and it was really helping, then I’d probably use it more and more.” – P4 (RH)
@@ 6.3. RQ2: Perceptions of Uncertainty
All participants in Conditions (R) and (RH) (P3–P5, P7) appreciated the presence of ranges. P5 compared the tool’s range to his own experiences: “I was always over the average. So, to me, I was kind of in my head using that as a low.” – P5 (R) For participants in the other conditions, P1 and P2 indicated that they would’ve preferred to have had ranges. However, P6 suggested that ranges may lead to disappointment when they are used for goal-setting: “If they earn less than [the higher number], then they probably might feel disappointed in the tool through no fault of its own, right? If you say \$12 to \$18, and it comes in at \$14, […] I could understand how folks might look at that as a let down.” – P6 (D) Thus, range-based uncertainty was useful for decision-making, but needed to be calibrated against expectations.
@@ 7.1. Key Findings and Implications
Interactivity could help to maintain trust over time. Losses of trust can also be mitigated retrospectively based on trust repair strategies (Pareek et al., 2024). Based on our qualitative findings, we hypothesise two mechanisms by which interactivity could help to maintain and repair trust. First, interactivity may enhance perceptions of control. The modes of interaction suggested by our participants, such as hourly breakdowns and question-answering, would assure users that the AI has the intent and agency to capture and learn from their preferences (Pareek et al., 2024). Second, interactivity could help users to better recall their experiences and decisions. Until the interviewer probed further, most interview participants could not recall whether their earnings significantly differed from the tool’s estimates.
@@ 7.2. Limitations
Our work has two primary limitations. First, we cannot claim that the design of our tool was optimal for engendering trust. Our focus was on testing how designs for exposing uncertainty would impact trust and reliance. Thus, we attempted to isolate the effect of this design dimension by refining the tool through a formative pilot study (Appendix B). Nevertheless, further improvements may have been possible. For instance, we could not provide retrospective breakdowns of participants’ earnings due to data availability limitations. Thus, design choices orthogonal to the exposure of uncertainty may have impacted participants’ trust and reliance.
Second, despite our best efforts, our sample of drivers was limited. These individuals were at least aware, if not active users, of the Gridwise app, and thus they may have been more focused on their outcomes than the general gig driver population. The trust of users in an AI decision aid is contingent upon their domain knowledge (Kim et al., 2023; Wysocki et al., 2023), and — as we demonstrate (Section 6.1) — the extent to which they integrate the decision aid into their existing routines. A future study aimed at a broader population of gig drivers could uncover additional insights by explicitly controlling for factors such as full-time status and driver tenure. We were also unable to reach participants who discontinued the user study. Future studies that follow up with such participants would be a valuable source of data on mechanisms of trust loss and repair in longitudinal settings.
@@ 7.3. Recommendations for Future Work
The paucity of similar longitudinal, in-situ studies in prior work is understandable given the logistical challenges we encountered. We stress the importance of observational studies to improve domain understanding as a basis for longitudinal interventional studies. Our pilot interviews helped us to design a tool that was task-aligned, which led participants to find value in it over repeated interactions. Furthermore, our study design aimed to increase perceived control while reducing user burden through flexibility in the scheduling of participation; customisability of the constraints on the AI decision aid; shorter survey instruments; and incremental compensation.
@@ Acknowledgements.
We thank Jenny T. Liang, Joshua Mateer, Jose M. del Alamo, Ian Yang, Chenyang Yang, Kaia Newman, Stephanie Milani, Niklas Kühl, and Nadia Nahar for their feedback and discussion on the study’s instruments, protocol, analysis, and manuscript. Also, we thank Linda Moreci for assistance with participant compensation, and Akshara Khare and Anagha Ravi Shankara for contributions to the recommendation tool’s implementation. This work was conducted in collaboration with Gridwise; we thank Kunal Prajapati, Karan Moudgil, Ryan Green, and Brian Finnamore. We were supported by NSF Grant CNS-1801316, a research grant from Mobility21, and the Tang Family Endowed Innovation Fund.
@@ A.4. Survey 4 — Post-Survey
    This item is an original question that prompts the participant to consider their interactions with other types of recommendation systems. It assesses the extent to which participants would appreciate increased exposure of uncertainty through range-based estimates in such systems.
@@ B.2. Methodology
To analyse these interviews, we used the same methodology as the longitudinal interviews (Section 6).
@@ B.3.1. Formative Questions
As the interviews were semi-structured, the script below focuses on the guiding questions that we asked participants. The interviewer also probed participants further depending on their responses.
@@ B.3.2. Evaluative Questions
  • Do you feel the recommended schedule gives you enough information to decide whether you would want to follow it?
@@ B.4. Participants
As with the longitudinal user studies (Section 4.1), participants were recruited from the user base of Gridwise in July 2023. Gridwise distributed recruitment messages to 500 users, but otherwise did not interact with participants. Recipients were sampled from Gridwise users in the United States who had completed at least one gig in a platform linked to the Gridwise app over the week preceding recruitment. We recruited 4 interview participants: